# Low-Altitude Wireless Networks: A Survey

### v1

@@ 5.2 Physical-Layer security
Physical layer security of LAWNs refers to protecting the confidentiality, integrity, and availability of information transmitted and processed within LAWNs by exploiting the inherent characteristics of the physical layer itself [129, 130]. This protection is achieved without relying solely on traditional cryptographic methods. By addressing the physical layer security concerns of the above issues, LAWNs can enhance resilience against sophisticated attacks, enable robust operation in contested environments, and provide lightweight security solutions suitable for resource-constrained UAV platforms, ultimately contributing to safer and more reliable UAV deployments. In general, physical layer security is commonly classified into two fundamental paradigms, namely the wiretap channel and covert communication. Currently, researchers have adopted beamforming techniques, artificial noise injection, cooperative jamming, power allocation strategies, and intelligent reflecting surfaces (IRS) to realize the physical layer security of LAWNs on sensing, communication, computing, and control aspects. For example, for physical layer security in integrated sensing and communication, [131] explored a system where a UAV transmits ISAC waveforms to securely communicate with IoT devices while simultaneously sensing the environment. This approach optimizes UAV trajectory and beamforming to maximize the average communication secrecy rate while ensuring accurate sensing capabilities. In the context of covert ISAC in LAWN systems, [132] studied covert ISAC for multi-antenna UAV systems using a unified waveform for communication and sensing, revealing a trade-off between covert communication and sensing accuracy by exploiting prior information of the adversary. Regarding physical layer security in UAV communication, [133] addressed wire-tap channel communications in 5G networks with eavesdroppers by proposing a virtual antenna array with collaborative beamforming. To enhance security, UAV positions and excitation current weights are optimized, aiming to improve total secrecy rates and reduce sidelobe levels through an improved multi-objective dragonfly algorithm. Furthermore, to satisfy the heterogeneous security requirements in low-altitude UAV systems, where some users demand secrecy while others require covertness, the authors in [134] investigated multi-user cooperative secret and covert communications for UAV networks. For the physical layer security aspect within computing, [135] investigated a secure UAV MEC system where ground users offload tasks to a legitimate UAV amidst eavesdropping UAVs. It enhances security by employing jamming signals from the legitimate UAV and users to maximize minimum secrecy capacity, jointly optimizing aspects like offloading ratio and computing capacity to balance security and latency. [136] focused on secure and energy-efficient UAV-MEC with RIS assistance, aiming to maximize minimum energy efficiency against an eavesdropper. The research ensured security through artificial noise transmission and optimizes computation tasks/CPU frequency allocation alongside other parameters, demonstrating strong performance even with an eavesdropper present. Focusing on physical layer security in control, [137] proposed a secret and covert transmission strategy for ground-to-UAV links using non-orthogonal multiple access (NOMA). This method allows simultaneous transmissions with different security needs, enabling an adversary to attempt both data eavesdropping and transmission detection. Security is rigorously ensured with optimization targeting maximized rates through beamforming and trajectory control. [138] improved covert communication range by using a UAV relay and finite blocklength transmissions, aiming to maximize secure bits against a warden. It achieves this by optimizing the warden’s threat potential and jointly controlling transmission parameters under error constraints, leveraging signal uncertainty at the warden for physical layer covertness.
@@ 5.3 Network-Layer security
For example, addressing network layer security challenges for UAV-based sensing, [139] designed an intrusion detection and response system to identify and mitigate attacks that compromise the integrity of sensed data, including false information dissemination and GPS spoofing. To ensure network layer security in UAV communication, [140] introduced self-adaptive intrusion detection (SID)-UAV, which employs a self-matching system to detect safe communication routes, coupled with a distributed architecture for investigation and response to malicious UAVs. The work of [141] enhanced network layer security for autonomous aerial vehicle (AAV)-aided short-packet transmissions against detection and eavesdropping by optimizing for maximum covert secrecy energy efficiency through a dual-loop iterative algorithm for user association, power control, and trajectory planning. Addressing network layer security in UAV edge computing, [142] presented a cyber-defense mechanism based on a non-cooperative game to protect against network and offloading threats. The solution aims to maximize protection while minimizing the computational and energy costs associated with the defense. In response to the data leakage risk in UAV-assisted traffic surveillance at the network layer, [143] utilized a probabilistic data structure approach for cyber-threat detection and efficient data management, minimizing computational overhead in the process. Regarding network layer security for control, [144] proposed using sliding mode control to enhance UAV security. The system dynamically reconfigures to adapt to network-level threats like GPS spoofing and jamming, ensuring robust control even under attack through anomaly detection and fault tolerance. [145] tackled network layer access control challenges in distributed UAV swarm control by using a blockchain-based scheme. It ensures secure authentication and data integrity against identity spoofing and data injection, employing dynamic key management for task and identity groups.
@@ 6.1 LAA Structuring
LAA is a fundamental element for the successful implementation of LAWNs, where the proper organization and structuring of airspace become crucial to ensure both operational efficiency and safety. Nevertheless, the corresponding configuration and structuring of LAA remain in the early stage. The performance and management efficiency of LAA will directly affect the operation of urban air mobility (UAM) systems, such that the LAA structure must not only accommodate uncrewed drones but also facilitate the integration with manned aircraft. Against this background, the SESAR program [146] defined a four-stage (U1-U4) development strategy for LAA. Specifically, the LAA corresponding to U1 is characterized by significant regionalization with clear no-fly zones and restricted areas, achieving only basic operational levels. The operational management system is formed in U2, improving the degree of openness and expanding the range of low-altitude aircraft, while the airspace in the U3 stage is further opened to enable simultaneous operation of high-density and multi-task missions. Finally, LAA is fully opened, which allows low-altitude aircraft and traditional airplanes to operate in the same airspace at intersecting altitudes. According to the ICAO framework [147], the entire airspace is divided into categories A, B, C, D, E, and G, with the G category being non-controlled airspace, while the others are controlled airspace. LAA primarily corresponds to categories G and E, but it may partially overlap with other categories in emergencies. The ultimate goal of LAA planning is how to allow for the joint operation of low-altitude aircraft and traditional aviation.
@@
In contrast to high-altitude airspace, LAA is significantly influenced by terrestrial environments, e.g., available flight altitudes, ground obstacles, aircraft noise, and weather conditions [148, 149]. Generally, these factors can be categorized into safety, environmental, and social concerns. Among them, safety is the most paramount factor in low-altitude operations with meteorological conditions and ground obstacles being particularly critical. For instance, adverse weather such as strong winds/rain, and low temperatures can reduce aircraft endurance, impair battery life, and increase collision risks. Although weather forecasting and pre-planned trajectories can mitigate some unfavorable impacts, real-time predictions remain challenging. Moreover. environmental factors like noise and visual pollution also play a significant role in public acceptance of LAWNs, especially in urban areas. The work of [150] showed that drone noise has become a key concern, as the public is more sensitive to drone noise than to vehicle noise. In parallel, visual pollution is another significant issue when dense drones fly overhead, affecting public emotions. On top of this, social factors, particularly in terms of privacy and equality preservation, influence public acceptance as well. UAVs’ ability to capture personal data during operations raises privacy concerns, and low-altitude services are often perceived as a privilege for high-income groups, leading to resistance from lower-income groups [151]. To address these challenges, existing research has structured LAA into five primary types based on various principles. i.e., pipeline airspace, corridor airspace, layered airspace, blocky airspace, and free airspace [152]. Each type presents unique design challenges and advantages, and various studies have explored their feasibility, limitations, and applications.
@@
1) Pipeline Airspace: As shown in Fig. 6(a), pipeline airspace is the most commonly used form of structured airspace, typically characterized by the establishment of flight path structures or virtual three-dimensional zones within the LAA, akin to road networks on the ground. The pipeline does not necessarily adhere to a fixed geometric shape, and its width is generally designed to accommodate only a single aircraft at a time. Aircraft within the pipeline operate in a unidirectional flight mode and can adjust their altitude, heading, and other parameters at designated intersection points. Notably, pipeline airspace offers a low degree of operational freedom but ensures high safety levels. For instance, the Metropolis project in Europe, which serves as a representative model of pipeline airspace, has designed independent pipelines for both uncrewed and manned aircraft [153]. The pipeline airspace consists of several key components: connector pipelines linking the ground to the lowest pipeline layer for takeoff and landing; horizontal pipelines connecting nodes within the same pipeline layer; and transition pipelines bridging different pipeline layers. When an aircraft passes through a marked node, it occupies that node for a specified period, with only one aircraft permitted to occupy the node at any given time [151]. Each node maintains an interval list to track the expected occupancy duration of other aircraft. Airspace resources are allocated based on the sequence in which aircraft enter the node, and new aircraft are only allowed to pass through once the node’s time slot becomes available. It is important to note that the number of marked nodes in each pipeline layer increases as the altitude decreases for safety.
@@
2) Corridor Airspace: Compared to pipeline airspace, corridor airspace offers a broader coverage with fewer establishment requirements. As observed in Fig. 6(b), it is typically organized in a regular geometric shape, characterized by two key components in terms of a centerline and a width for meeting certain performance criteria. The corridor airspace allows multiple aircraft to fly in parallel following a bidirectional flight pattern. Moreover, corridors can be designed as single or multi-layered airspace to meet traffic flow requirements to increase LAWN capacity. Corridor airspace offers significant advantages such as fewer flight restrictions and greater resilience, where the horizontal and vertical separations among drones are ensured, allowing for maximum flight freedom under safe operational conditions [154]. To further unlock the potential of corridor airspace, new configurations are investigated to accommodate higher traffic demands. Embraer-X, a Brazilian aerospace company, has proposed dividing LAA into corridors, road segments, and boundary points. In particular, the corridors are dynamically determined by GPS waypoints to fulfill various air traffic flow requirements and improve system efficiency [155]. The authors in [156] introduced multi-layer air corridors, divided into top, middle, and bottom layers, in which the top and bottom layers are designated for drone flying purposes along north-south and east-west directions, respectively. While the middle layer serves as a transitional area for aircraft to temporarily wait and change directions, and thus effectively reducing the collision risk.
@@
3) Layered Airspace: Layered airspace refers to the division of LAA into several separate layers, each with the same or varying vertical intervals. Since layered airspace is primarily divided in the vertical direction, the operational rules of aircraft in each layer can be effectively controlled, and aircraft operating at different altitudes will not interfere with each other. It is noteworthy that this structured design enhances airspace safety and lowers the technical requirements at the expense of freedom of airspace utilization degradation. In response to the drone market requirements, Amazon proposed a layered LAA design for lightweight drones operating below 500 ft [157]. In particular, the framework consists of a low-speed traffic layer (below 200 ft) for non-transport missions or UAVs with limited onboard capabilities, a high-speed transport layer (200-400 ft) for faster UAVs capable of autonomous navigation and conflict avoidance in complex environments, and a no-fly zone (400-500 ft) restricted to emergency use, as depicted in Fig. 6(c).

### v2

@@ 5.2 Physical-Layer security
Physical layer security of LAWNs refers to protecting the confidentiality, integrity, and availability of information transmitted and processed within LAWNs by exploiting the inherent characteristics of the physical layer itself [wang2021physical, 10168298]. This protection is achieved without relying solely on traditional cryptographic methods. By addressing the physical layer security concerns of the above issues, LAWNs can enhance resilience against sophisticated attacks, enable robust operation in contested environments, and provide lightweight security solutions suitable for resource-constrained UAV platforms, ultimately contributing to safer and more reliable UAV deployments. In general, physical layer security is commonly classified into two fundamental paradigms, namely the wiretap channel and covert communication. Currently, researchers have adopted beamforming techniques, artificial noise injection, cooperative jamming, power allocation strategies, and intelligent reflecting surfaces (IRS) to realize the physical layer security of LAWNs on sensing, communication, computing, and control aspects. For example, for physical layer security in integrated sensing and communication, [xiu2024improving] explored a system where a UAV transmits ISAC waveforms to securely communicate with IoT devices while simultaneously sensing the environment. This approach optimizes UAV trajectory and beamforming to maximize the average communication secrecy rate while ensuring accurate sensing capabilities. In the context of covert ISAC in LAWN systems, [xx] studied covert ISAC for multi-antenna UAV systems using a unified waveform for communication and sensing, revealing a trade-off between covert communication and sensing accuracy by exploiting prior information of the adversary. Regarding physical layer security in UAV communication, [li2021physical] addressed wire-tap channel communications in 5G networks with eavesdroppers by proposing a virtual antenna array with collaborative beamforming. To enhance security, UAV positions and excitation current weights are optimized, aiming to improve total secrecy rates and reduce sidelobe levels through an improved multi-objective dragonfly algorithm. Furthermore, to satisfy the heterogeneous security requirements in low-altitude UAV systems, where some users demand secrecy while others require covertness, the authors in [10681761] investigated multi-user cooperative secret and covert communications for UAV networks. For the physical layer security aspect within computing, [zhou2019secure] investigated a secure UAV MEC system where ground users offload tasks to a legitimate UAV amidst eavesdropping UAVs. It enhances security by employing jamming signals from the legitimate UAV and users to maximize minimum secrecy capacity, jointly optimizing aspects like offloading ratio and computing capacity to balance security and latency. [nasir2024secure] focused on secure and energy-efficient UAV-MEC with RIS assistance, aiming to maximize minimum energy efficiency against an eavesdropper. The research ensured security through artificial noise transmission and optimizes computation tasks/CPU frequency allocation alongside other parameters, demonstrating strong performance even with an eavesdropper present. Focusing on physical layer security in control, [xu2025collaborative] proposed a secret and covert transmission strategy for ground-to-UAV links using non-orthogonal multiple access (NOMA). This method allows simultaneous transmissions with different security needs, enabling an adversary to attempt both data eavesdropping and transmission detection. Security is rigorously ensured with optimization targeting maximized rates through beamforming and trajectory control. [chen2021uav] improved covert communication range by using a UAV relay and finite blocklength transmissions, aiming to maximize secure bits against a warden. It achieves this by optimizing the warden’s threat potential and jointly controlling transmission parameters under error constraints, leveraging signal uncertainty at the warden for physical layer covertness.
@@ 5.3 Network-Layer security
For example, addressing network layer security challenges for UAV-based sensing, [sedjelmaci2017hierarchical] designed an intrusion detection and response system to identify and mitigate attacks that compromise the integrity of sensed data, including false information dissemination and GPS spoofing. To ensure network layer security in UAV communication, [fotohi2022self] introduced self-adaptive intrusion detection (SID)-UAV, which employs a self-matching system to detect safe communication routes, coupled with a distributed architecture for investigation and response to malicious UAVs. The work of [mao2024energy] enhanced network layer security for autonomous aerial vehicle (AAV)-aided short-packet transmissions against detection and eavesdropping by optimizing for maximum covert secrecy energy efficiency through a dual-loop iterative algorithm for user association, power control, and trajectory planning. Addressing network layer security in UAV edge computing, [sedjelmaci2019efficient] presented a cyber-defense mechanism based on a non-cooperative game to protect against network and offloading threats. The solution aims to maximize protection while minimizing the computational and energy costs associated with the defense. In response to the data leakage risk in UAV-assisted traffic surveillance at the network layer, [garg2018uav] utilized a probabilistic data structure approach for cyber-threat detection and efficient data management, minimizing computational overhead in the process. Regarding network layer security for control, [ale2024advanced] proposed using sliding mode control to enhance UAV security. The system dynamically reconfigures to adapt to network-level threats like GPS spoofing and jamming, ensuring robust control even under attack through anomaly detection and fault tolerance. [xie2025blockchain] tackled network layer access control challenges in distributed UAV swarm control by using a blockchain-based scheme. It ensures secure authentication and data integrity against identity spoofing and data injection, employing dynamic key management for task and identity groups.
6 Advanced AI-driven LAWNs
Traditional model-based optimization and control frameworks, while effective under static and predictable conditions, struggle to cope with the high-dimensional, dynamic, and uncertain nature of low-altitude environments. In this section, we delve into the integration of advanced AI technologies into LAWNs, focusing on the deployment of computationally efficient large AI models and agentic AI. In particular, large AI models enhance cognitive capabilities for complex decision-making and environmental understanding, and agentic AI empowers LAWNs to be autonomous and collaborate in dynamic environments, improving the scalability and resilience.
6.1 Large AI Model-based LAWNs
Large AI models, characterized by their extensive parameter dimensions and training on massive datasets, have made notable advancements across various domains, including natural language processing, computer vision, and autonomous systems. Prominent models, e.g., generative pre-trained transformers (GPT), bidirectional encoder representations from transformers (BERT), and similar architectures, possess the ability to learn intricate patterns and generalize across a wide range of tasks. This makes them particularly suitable for dynamic and complex environments. Hence, the integration of large AI models into LAWNs holds substantial potential for enhancing the cognitive capabilities of UAVs, facilitating real-time decision-making in unpredictable environments. A common approach to this integration involves utilizing large language models (LLMs) in physical layer designs, where pre-trained models are deployed at both the UAVs and controller terminal to enable effective online inference. For example, the work of [Xiao2024] illustrates how large AI models can be employed in UAV path planning, where they enable the UAVs to dynamically adjust their flight paths based on real-time environmental data, such as wind and obstacle conditions. Similarly, the work of [Javaid2024] explores how large language models (LLMs) can enhance UAV communication by enabling more intuitive interactions between UAVs and ground control stations. This capability allows for more adaptive and flexible communication, improving operational coordination, especially in complex scenarios like disaster relief or surveillance operations.
Despite their impressive capabilities, deploying large AI models in LAWNs presents significant challenges, primarily due to the substantial computational and memory requirements of these models. The increasing size of model parameters, often reaching billions, imposes a heavy computational burden, which exceeds the resources available in many UAVs and edge devices. The need for real-time decision-making further complicates the deployment of such models, as low latency is critical in many applications. To circumvent these challenges, two prominent distributed strategies have been proposed, i.e., model distribution and data distribution. Both approaches aim to alleviate the computational load on individual UAVs and edge devices by distributing the processing requirements across a network of devices.
Model distribution involves partitioning a large AI model across multiple devices or computational nodes, such as UAVs or edge servers. Instead of running the entire model on a single device, each UAV or edge server processes a portion of the model. This distributed approach allows for parallel processing, which significantly reduces the computational burden on each individual device. The work of [Piggott2023] demonstrated how model distribution can be effectively implemented in UAV networks, where large AI models are divided across several UAVs, enabling real-time inference without overwhelming the resources of any single UAV. This technique enhances the scalability of large models, making it possible to deploy them in a distributed manner without compromising system performance. Data distribution, on the other hand, focuses on distributing the data processing tasks across multiple devices to minimize the need for large-scale data transmission. In the context of LAWNs, this approach ensures that data collected by UAVs is processed locally, reducing the communication overhead and enabling real-time decision-making. The work of [Javaid2024] highlighted the use of federated learning in UAV networks, a data distribution method that allows multiple UAVs to collaboratively train a shared model by exchanging model updates rather than raw data. This approach not only preserves data privacy but also reduces the bandwidth required for communication between UAVs and ground control systems, making it particularly beneficial in applications where security and privacy are paramount.
However, it is important to note that both model distribution and data distribution entail substantial resource-intensive fine-tuning processes [10415235]. Therefore, high-dimensional computing (HDC) has attracted increasing attention as a computationally efficient alternative [chang2023recent]. HDC is a brain-inspired computing paradigm in which information is encoded into extremely high-dimensional vectors, often containing tens of thousands of components. Such representations naturally support large-scale parallelism and enable rapid processing, while exhibiting strong robustness against noise, distortions, and missing data. A fundamental property of these high-dimensional vectors is that independently generated vectors are almost orthogonal to one another. This behavior results from the randomization procedures during vector construction, which ensure that distinct vectors share very limited overlap [basaklar2021hypervector]. Consequently, meaningful representations tend to have large Hamming distances and low cosine similarity, providing substantial capacity for encoding diverse information with minimal interference. These characteristics enable HDC to rely on simple arithmetic operations, including addition for combining multiple vectors and multiplication for binding variables with their associated values [kanerva2009hyperdimensional]. Similarity measures, such as Hamming distance and cosine similarity, are used to determine how closely two vectors align, supporting accurate retrieval and comparison. Since these operations do not require heavy matrix computations, they are highly efficient and particularly suitable for UAVs operating in LAWNs, where memory, computation capability, and energy supply are inherently limited.
HDC has also shown strong potential for improving wireless communication reliability in scenarios relevant to LAWNs. Its high-dimensional representations can tolerate interference and signal distortions, allowing communication to remain stable even when multiple devices transmit simultaneously. A representative example is the direct transmission of high-dimensional vectors from a group of sensing devices. Instead of coordinating access or assigning dedicated channels, each device simply sends the vector that corresponds to its measurement. The receiver then observes the superimposed vector and determines the underlying sensing events by comparing the received vector with a set of reference patterns. By doing so, HDC enables efficient UAV reporting without traditional access control. For example, the authors in [jakimovski2012collective] considered a scenario where the receiver detects temperature conditions and determines how many devices experienced a particular temperature by analyzing the combined vector. Furthermore, the study in [kleyko2012dependable] generated composite vectors through binding and superposition to encode local sensing information at each device. The receiver successfully interpreted the embedded information using predefined base vectors, demonstrating that HDC can maintain reliable information extraction even in dense and interference-prone environments. These results indicate that HDC offers a practical and lightweight computational method for LAWNs, enabling robust communication and efficient information exchange while accommodating the limitations of UAV platforms.
6.2 Agentic AI-based LAWNs
Although lightweight large AI models substantially strengthen LAWNs by offering advanced perception and inference capabilities, the operational role of these models in current designs remains largely passive. In most architectures, large models are tied to predefined objectives, rely on offline pretraining, and function primarily as static function approximators that map inputs to outputs under fixed task specifications[Javaid2024]. The behavior is effectively constrained by the training data distribution and objective design, which limits the ability to autonomously revise goals, reinterpret evolving mission contexts, or sustain long-horizon coordination with other agents. Such rigidity is particularly problematic in LAWNs, where UAVs operate in nonstationary environments, face time-varying and sometimes conflicting mission objectives, and must continuously interact with other aircraft and ground infrastructure.
Agentic AI offers a promising paradigm to overcome these limitations by elevating LAWNs from passive optimization engines to active, goal-directed autonomous systems. By embedding reasoning, memory, tool utilization, and self-evolution into network entities, agentic AI enables UAVs and ground infrastructure to adapt trajectories, allocate resources, and collaborate in swarms in a mission-aligned manner [sapkota2025uavs, gao2025agentic]. Through continual learning and shared intent representations, agentic LAWN systems are capable of mitigating performance drift induced by mobility and partial observability, thereby sustaining resilient closed-loop operation in complex low-altitude scenarios [acharya2025agentic]. In this context, large models and learning modules are no longer treated as static components; instead, they form the core of deliberative agents that can interpret goals, plan actions, and revise strategies online as the environment and mission requirements evolve.
It is worth mentioning that, within the agentic paradigm, task integration emerges as the central organizing principle for AI-based LAWNs. Rather than treating sensing, communication, control, and cognition as loosely coupled functions, the network is structured around mission-level tasks that are decomposed into several interdependent subtasks, which are allocated to heterogeneous agents and continuously re-evaluated as environments change. In particular, a task-oriented cognitive layer can be adopted to interpret high-level intent and fuse multi-modal observations, while lower layers implement coordinated functions, e.g., ISAC, wireless networked control, and edge intelligence. By doing so, UAVs and ground nodes can reason jointly about trade-offs in communication reliability, control precision, and sensing accuracy to adapt behavior based on both local context and global performance metrics [feng2025multi, zhou2024comprehensive]. In this way, the mission goal is translated into consistent cross-layer decisions, followed by gradually reshaping task priorities and resource allocations based on real-time feedback from execution.
Building on this integrated view, some recent research has begun to explore agentic optimization solutions that couple mathematical coordination frameworks with autonomous decision-making. Optimization-based approaches, originally formulated in a centralized manner, are extended to distributed and agent-based schemes in which each UAV acts as a decision-making agent handling local objectives and constraints while obeying global mission requirements. For example, distributed optimization has been employed to jointly design multi-role UAV functionality switching and trajectories for secure task offloading in UAV-assisted mobile edge computing, where local agents iteratively coordinate to satisfy communication and computation constraints under dynamic conditions [zhong2024distributed]. In parallel, LAWN-wide closed-loop designs have been formulated that integrate sensing, communication, and control into unified optimization problems, demonstrating that jointly tuned resource allocation and controller design can significantly enhance positioning accuracy and service reliability compared with decoupled baselines [wang2024semantic, zhou2024comprehensive]. Complementing optimization-driven coordination, agentic AI solutions build learning-based autonomy on top of this substrate. At the single-agent level, reinforcement learning (RL) has been widely adopted to endow individual UAVs with adaptive decision-making capabilities in complex, uncertain environments. Typical implementations rely on actor–critic architectures, deep Q-networks, or proximal policy optimization to learn control and path-planning policies directly from interaction with the environment [kaelbling1996reinforcement, konda1999actor, schulman2017proximal]. In soft actor–critic–based designs, for instance, a UAV agent observes its flight state and local environment and outputs continuous control actions that track reference trajectories while avoiding obstacles and following energy and communication constraints [cheng2020autonomous]. The entropy-regularized objective promotes a balance between exploitation of known good behaviors and exploration of new strategies, improving robustness to modeling errors and unforeseen disturbances. From the LAWN perspective, such a controller functions as a local decision engine that interprets mission-level commands and translates them into fine actuation in real time.
When multiple UAVs and ground nodes need to coordinate, multi-agent reinforcement learning (MARL) extends this paradigm to networked settings, where centralized training and decentralized execution are applied[kopic2024collaborative]. Specifically, each UAV is modeled as an agent with partial local observations and a policy that must adapt under the nonstationary dynamics induced by other learning agents. To name a few, the authors in [cui2019multi] investigated MARL-based resource allocation schemes that enable UAVs to share spectral and power resources while jointly maximizing network throughput and fairness. Furthermore, the work of [liang2023multi] developed a multi-UAV collaborative search and attack framework that formulated exploration, target engagement, and collision avoidance in unknown environments as a multi-agent decision problem. In fact, such learned coordination strategies are intractable to engineer manually, highlighting the value of agentic learning in realizing cooperative, mission-aware behaviors in LAWNs, even though issues such as nonstationarity, credit assignment, and training stability remain challenging.
A further step toward cognitively richer agentic AI integrates LLMs as reasoning engines for UAV agents. For example, LLM-based UAV path-planning frameworks have been proposed for industrial scenarios, where an LLM is employed to interpret mission descriptions, query external tools or simulators, and generate high-level plans that are subsequently translated into executable trajectories for individual UAVs [wang2024survey]. In addition, the studies in [Javaid2024, Li2025] provided a systematic overview of how LLMs can support intent understanding, dialogue-based supervision, and explainable decision-making in LAWN systems. Operational prototypes such as Net-GPT further demonstrated LLM-empowered intermediaries that maintain contextual information, enforce safety policies, and coordinate interactions among UAVs, human operators, and external services in real time [Piggott2023]. Collectively, these developments point to a transition toward cognitive agentic LAWNs, where LLMs handle high-level reasoning, task decomposition, and tool orchestration, while specialized learning or optimization modules are responsible for low-level control and communication under explicit safety and regulatory constraints.
@@ 7.1 LAA Structuring
Figure 7: The supported various LAWN services associated with different U-space stages based on [juntong2025low].
@@
LAA is a fundamental element for the successful implementation of LAWNs, where the proper organization and structuring of airspace become crucial to ensure both operational efficiency and safety[jia2025hierarchical, 11159633]. Nevertheless, the corresponding configuration and structuring of LAA remain in the early stage. The performance and management efficiency of LAA will directly affect the operation of urban air mobility (UAM) systems, such that the LAA structure must not only accommodate uncrewed drones but also facilitate the integration with manned aircraft. According to the ICAO framework [FAA2022], the entire airspace is divided into categories A, B, C, D, E, and G, with the G category being non-controlled airspace, while the others are controlled airspace. LAA primarily corresponds to categories G and E, but it may partially overlap with other categories in emergencies. The ultimate goal of LAA planning is to allow for the joint operation of low-altitude aircraft and traditional aviation. Against this background, the SESAR program [SESAR2017] defined a four-stage (U1-U4) development strategy for LAA, which is particularized as:
  • U1 corresponds to the initial stage of U-space deployment, defined in 2019 for class G airspace below 400 ft and aimed at supporting mixed crewed and uncrewed operations under relatively low traffic density. At this level, operations are typically confined to restricted or predesignated areas, and the service set remains intentionally limited. The key functions consist of basic registration of unmanned aircraft and geo-awareness capabilities, enabling operators to understand and respect airspace restrictions while laying the groundwork for more advanced U-space services in subsequent phases.
  • U2 builds on this foundation by enabling beyond visual LoS (BVLoS) operations in higher-density environments, extending into portions of controlled airspace. In this phase, U-space introduces a procedural interface with conventional air traffic control (ATC), allowing uncrewed operations to be coordinated with existing ATM procedures. The service portfolio expands to drone operation management functions, including flight planning and approval, real-time tracking of flights, and the provision of dynamic airspace information, thereby supporting safer and more predictable BVLoS activities.
  • U3 represents a further step toward advanced U-space services, targeting BVLoS operations in dense and complex airspace such as suburban areas and emerging urban air mobility corridors. This stage assumes more sophisticated and higher levels of automation in the interaction between U-space and ATC. Service capabilities are enhanced to include advanced geo-awareness, dynamic geofencing, dynamic capacity management, and conflict detection and resolution functions. According to current planning, U3-level services are expected to reach operational maturity around 2027.
  • U4 envisions comprehensive integration of unmanned aircraft into very dense urban airspace, with a correspondingly high level of automation for both vehicles and the supporting network and U-space infrastructure. At this stage, U-space and traditional ATM services are expected to be fully implemented and tightly coupled, enabling seamless management of crewed and uncrewed traffic in highly demanding environments. Full realization of U4 capabilities is anticipated around 2035 or later.
@@
Along with this evolution, the service sets associated with U1 through U4 are summarized in Fig. 7. In contrast to high-altitude airspace, LAA is significantly influenced by terrestrial environments, e.g., available flight altitudes, ground obstacles, aircraft noise, and weather conditions [Bauranov2021, Tang2022]. Generally, these factors can be categorized into safety, environmental, and social concerns. Among them, safety is the most paramount factor in low-altitude operations with meteorological conditions and ground obstacles being particularly critical. For instance, adverse weather such as strong winds/rain, and low temperatures can reduce aircraft endurance, impair battery life, and increase collision risks. Although weather forecasting and pre-planned trajectories can mitigate some unfavorable impacts, real-time predictions remain challenging. Moreover, environmental factors like noise and visual pollution also play a significant role in public acceptance of LAWNs, especially in urban areas. The work of [Schaffer2021] showed that drone noise has become a key concern, as the public is more sensitive to drone noise than to vehicle noise. In parallel, visual pollution is another significant issue when dense drones fly overhead, affecting public emotions. On top of this, social factors, particularly in terms of privacy and equality preservation, influence public acceptance as well. UAVs’ ability to capture personal data during operations raises privacy concerns, and low-altitude services are often perceived as a privilege for high-income groups, leading to resistance from lower-income groups [Sunil2015]. To address these challenges, existing research has structured LAA into five primary types based on various principles. i.e., pipeline airspace, corridor airspace, layered airspace, blocky airspace, and free airspace [juntong2025low]. Each type presents unique design challenges and advantages, and various studies have explored their feasibility, limitations, and applications.
@@
1) Pipeline Airspace: As shown in Fig. 8(a), pipeline airspace is the most commonly used form of structured airspace, typically characterized by the establishment of flight path structures or virtual three-dimensional zones within the LAA, akin to road networks on the ground. The pipeline does not necessarily adhere to a fixed geometric shape, and its width is generally designed to accommodate only a single aircraft at a time. Aircraft within the pipeline operate in a unidirectional flight mode and can adjust their altitude, heading, and other parameters at designated intersection points. Notably, pipeline airspace offers a low degree of operational freedom but ensures high safety levels. For instance, the Metropolis project in Europe, which serves as a representative model of pipeline airspace, has designed independent pipelines for both uncrewed and manned aircraft [metropolis2014]. The pipeline airspace consists of several key components: connector pipelines linking the ground to the lowest pipeline layer for takeoff and landing; horizontal pipelines connecting nodes within the same pipeline layer; and transition pipelines bridging different pipeline layers. When an aircraft passes through a marked node, it occupies that node for a specified period, with only one aircraft permitted to occupy the node at any given time [Sunil2015]. Each node maintains an interval list to track the expected occupancy duration of other aircraft. Airspace resources are allocated based on the sequence in which aircraft enter the node, and new aircraft are only allowed to pass through once the node’s time slot becomes available. It is important to note that the number of marked nodes in each pipeline layer increases as the altitude decreases for safety.
@@
2) Corridor Airspace: Compared to pipeline airspace, corridor airspace offers a broader coverage with fewer establishment requirements. As observed in Fig. 8(b), it is typically organized in a regular geometric shape, characterized by two key components in terms of a centerline and a width for meeting certain performance criteria. The corridor airspace allows multiple aircraft to fly in parallel following a bidirectional flight pattern. Moreover, corridors can be designed as single or multi-layered airspace to meet traffic flow requirements to increase LAWN capacity. Corridor airspace offers significant advantages such as fewer flight restrictions and greater resilience, where the horizontal and vertical separations among drones are ensured, allowing for maximum flight freedom under safe operational conditions [airbus2018]. To further unlock the potential of corridor airspace, new configurations are investigated to accommodate higher traffic demands. Embraer-X, a Brazilian aerospace company, has proposed dividing LAA into corridors, road segments, and boundary points. In particular, the corridors are dynamically determined by GPS waypoints to fulfill various air traffic flow requirements and improve system efficiency [embraerx2019]. The authors in [muna2021] introduced multi-layer air corridors, divided into top, middle, and bottom layers, in which the top and bottom layers are designated for drone flying purposes along north-south and east-west directions, respectively. While the middle layer serves as a transitional area for aircraft to temporarily wait and change directions, and thus effectively reducing the collision risk.
@@
3) Layered Airspace: Layered airspace refers to the division of LAA into several separate layers, each with the same or varying vertical intervals. Since layered airspace is primarily divided in the vertical direction, the operational rules of aircraft in each layer can be effectively controlled, and aircraft operating at different altitudes will not interfere with each other. It is noteworthy that this structured design enhances airspace safety and lowers the technical requirements at the expense of freedom of airspace utilization degradation. In response to the drone market requirements, Amazon proposed a layered LAA design for lightweight drones operating below 500 ft [amazon2015]. In particular, the framework consists of a low-speed traffic layer (below 200 ft) for non-transport missions or UAVs with limited onboard capabilities, a high-speed transport layer (200-400 ft) for faster UAVs capable of autonomous navigation and conflict avoidance in complex environments, and a no-fly zone (400-500 ft) restricted to emergency use, as depicted in Fig. 8(c).